# Crepe: A Mobile Screen Data Collector Using Graph Query

### v1

@@ Abstract.
Collecting mobile datasets remains challenging for academic researchers due to limited data access and technical barriers. Commercial organizations often possess exclusive access to mobile data, leading to a “data monopoly” that restricts the independence of academic research. Existing open-source mobile data collection frameworks primarily focus on mobile sensing data rather than screen content, which is crucial for various research studies. We present Crepe, a no-code Android app that enables researchers to collect information displayed on screen through simple demonstrations of target data. Crepe utilizes a novel Graph Query technique which augments the structures of mobile UI screens to support flexible identification, location, and collection of specific data pieces. The tool emphasizes participants’ privacy and agency by providing full transparency over collected data and allowing easy opt-out. We designed and built Crepe for research purposes only and in scenarios where researchers obtain explicit consent from participants. Code for Crepe will be open-sourced to support future academic research data collection.
journalyear: 2023copyright: rightsretainedjournal: PACMHCIjournalyear: 2024journalvolume: 6journalnumber: UISTarticle: 1publicationmonth: 11doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Datasets collected from apps running on mobile devices are valuable resources in Human Computer Interaction (HCI) and social sciences (Raento et al., 2009), but collecting mobile app datasets has remained challenging for academic researchers. In practice, access to data siloed in mobile apps (Chandramouli et al., 2015) is often exclusively possessed by the smartphone platforms (e.g. iOS, Google Android) or individual products’ developers (Generosi et al., 2020). Academic researchers have to either use public developer APIs (Trupthi et al., 2017) or directly collaborate with products’ creators (Kramer et al., 2014) to obtain mobile datasets. However, many researchers have limited resources and technical expertise to build their own data collectors. Meanwhile, the availability of developer APIs are subject to corporations’ discretionary policies (Reddit, 2023; Barnes, 2023). Also, when researchers collaborate with commercial organizations to access mobile data (Kramer et al., 2014; Deka et al., 2017), they often need to align with the organizations’ business goals and public images, restricting the independence of academic research (Kramer et al., 2014). These phenomena are recognized as commercial organizations’ “data monopoly” in previous work (Jia-Jun Li et al., 2022) and pose challenges for empirical research related to mobile devices.
@@
Many open-source mobile data collection frameworks have been created to support academics; however, most mature solutions focus on mobile sensing data (Ferreira et al., 2015; Aharony et al., 2011; Cardone et al., 2013; Tangmunarunkit et al., 2015; Schueller et al., 2014; Ferreira et al., 2015; Andone et al., 2016; Cao et al., 2015) instead of screen information collection. Screen information i.e., the various data displayed on mobile user interface (UI) screens, is an important data source to enable various research studies (Thomas et al., 2020; Robinson et al., 2017; Supanitayanon et al., 2020; Hill and Monroy-Hernández, 2017). A major technical challenge is the difficulty to reliably and automatically identify which user interface screens contain the target data, and then locate and collect it. As a result, many mobile screen collectors require users to actively input the data or upload screenshots containing the target data (Calacci and Pentland, 2022). Some collectors continuously record user screens to get the target information, which is less efficient and poses privacy concerns (Krieter, 2019).
In this work, our goal is to build a low-code tool to support the collection of screen information and user interaction data on mobile devices. To tackle the above challenges in mobile data collection, we designed and implemented a novel query language, Graph Query, which augments mobile UI screen structures to support the flexible identification, location, and collection of specific data pieces. On top of Graph Query, we designed and built an Android collector app named Crepe [^1] Crepe stands for “Collector for Research Experiments of Participant Experiences”.. Crepe supports researchers to collect screen data through simple, no-code demonstrations of the target data to collect, as well as easy, scalable sharing of collectors with participants.
We designed Crepe with a special focus on data collection participants’ privacy. The participants can download Crepe and easily join a data collection study by entering a collector ID shared by a researcher. Users will have full transparency over what data is collected from their phones and can leave any data collection at any time. We intend Crepe to be used for research purposes only, and require researchers to obtain explicit consent from participants and fully disclose the data being collected to participants.
@@
    A novel query language to effectively identify, locate, and collect target data on UI screens;
@@
    The first Android data collector app, Crepe, to support flexible, customizable, low-code UI screen data collection using a programming by demonstration (PBD) paradigm (Cypher and Halbert, 1993);
@@
    A user study with 7 participants in a realistic data collection task that uncovered the real-world performance, limitations, and future development directions of Crepe.
@@ 2.1. Understanding Human Activities via Mobile Data
Mobile devices, especially smartphones, are now ubiquitous in everyday lives. Mobile applications harness various information about user social activities, which are embedded in users’ interactions, app usage patterns, and preferences. As a result, researchers in HCI and social science have been leveraging rich data gained from mobile apps to better improve user experiences (Deka et al., 2017; Wang et al., 2023), uncover societal trends (Vieweg et al., 2010; Starbird et al., 2014; Starbird and Palen, 2010), and understanding user behaviors and attitudes (Fakhradyan, 2021; De et al., 2017; Balaban et al., 2022; Kramer et al., 2014).
@@
2.1.1. Utilizing Behavioral Data Stream for Analysis
Researchers from a diverse background, such as social science or business, mine mobile data for behavioral patterns and insights. Experience Sampling Method (ESM) is a research methodology to obtain in situ data for constructing an understanding of subjects’ daily behaviors, feelings, and thoughts of participants (van Berkel et al., 2017; Yue et al., 2014). In the survey conducted by Berkel et al., smartphone-based ESM benefits researchers in improving data quality through validation, collecting rich multimodal data for context reconstruction, and enabling real-time data collection and analysis (van Berkel et al., 2017). ESM can provide social scientists with combined human and sensor data in addition to background logging (Raento et al., 2009). Aside from sensor data, ESM questionnaires can also capture external aspects of experience, such as time, place, activities, and companion (Hektner et al., 2007). For instance, Yue et al. adopted ESM to study the function of photos. They found that photos can be used to trigger memory during follow-up interviews and as a beneficial component in data analysis (Yue et al., 2014). In this work, our goal is to create a low-code mobile UI screen data collection tool. Our app Crepe can facilitate academic researchers to instrument screen data collections without the need to extensively develop customized data collectors, or have public access to data APIs.
@@ 3. Proposed Method
Our goal for the Crepe data collector is three-fold. First, we seek to empower researchers lacking programming expertise to easily instrument data collections. Second, we prioritize the privacy and agency of individual data collection participants. Lastly, we also want to address the following observed technical challenges in reliable mobile screen data collection, which hinder even researchers with technical skills:
@@
    The difficulty in flexibly and automatically detecting the appearance of target data on the screen;
@@
    The lack of precision in reliably pinpointing the location of target data;
@@
    The limited generalizability to diverse data types, especially for dynamically evolving data.
To get around these challenges, previous collectors have asked participants to either directly input target data or upload screenshots or screen recordings containing target data  (Calacci and Pentland, 2022; Krieter, 2019). However, these practices significantly increase participants’ efforts and reduce the reliability of the collected data’s quality.
@@ 3.1. The Crepe Data Collector User Experience
Figure 2. The workflow of Crepe for our main user groups: data collection researchers and participants. Researchers create a new data collector by demonstration (A) and share the collector ID with participants (B). Participants add the collector to their own devices (C), which runs in the background to collect the specified target data (D). The collected data is transmitted to a database (E) for the researcher to analyze and gain insights (F). The colors in the background indicate each user group’s experience involved in the holistic Crepe pipeline.
@@
The figure illustrates the workflow of the Crepe system with 6 labeled steps: A) The researcher creates a new data collector by specifying the time period, app package, data field and description through a graphical interface on their device. B) The researcher shares the collector ID with participants. C) Participants add the collector to their own devices by entering the shared collector ID into the Instagram app running on their phone. The collector details like start/end time are shown. D) On the participant’s device, the Crepe app automatically identifies and locates the target UI data elements (like ”walmart”, ”deucelove”) using Graph Query, based on the UI screenshot. E) The target data elements collected by Crepe on participant devices get transmitted to a database. F) The researcher analyzes the collected data from the database to gain insights for their study. Arrows show the flow of information between the researcher device, participant devices, and the data collection database.
@@
Researchers have two main tasks when using Crepe (Figure 2): creating a collector (A), and sharing collectors with study participants (B). A researcher can use Crepe’s graphical user interfaces to define the start and end dates of a collector, as well as the target app to collect data from.
@@
The most important yet trickiest step in defining a data collector is specifying the target data on screen to collect. In Crepe, our design goal is to minimize the complexity of this specification process for researchers, by only asking them to tap on the screen data to collect in the target app, and describe their intention to collect these data (Figure 2 (A)). Crepe automatically translate these two pieces of information into a formal, executable Graph Query, which includes information about the target data’s parent app package, component type, and relations with other UI entities on screen. This Graph Query is the core of a data collector and will be used to reliably identify and precisely locate target data on the study participants’ devices. This specification process follows a programming-by-demonstration paradigm (Cypher and Halbert, 1993) and provides a simple, no-code experience for researchers to specify the target data. (The technical detail will be further explained in the following subsection on Graph Query.) After the target data are specified and descriptions are added, a collector is created. Afterward, the researcher can copy and share the unique ID of this collector with participants (B). The collector will run on the participants’ device for the specified data collection periods. Afterwards, the researcher can receive the data collected in a self-defined database (E) and analyze it to uncover research insights (F). Our initial implementation of Crepe transmits collected data to a pre-configured Google Firebase Realtime Database that encrypts data both in transit and at rest [^3] https://firebase.google.com/support/privacy. We plan to support researchers’ customized configuration of their own database servers in the future.
@@
For study participants, they can easily add a collector to Crepe by entering the unique collector ID shared by the study’s researchers (C), using Crepe’s graphical user interfaces. After granting Crepe with the necessary Android Accessibility Service permission, the added collector(s) will start automatically running in the background of the participants’ devices (D). Guided by the graph queries associated with the participants’ added collectors, Crepe will selectively seek the appearance of target data exclusively when the designated target apps are active. Once the target data is detected, Crepe will scrape and transmit the relevant information to the researchers (E). The participant will be able to monitor the active status of the collector(s) they have joined in Crepe. To drop out of a study, participants can simply delete a collector from their Crepe app, and the collector will no longer run in the background.
Privacy considerations
@@
Protecting the privacy and agency of study participants is critical in our design of Crepe. We ensure individual participants have the full control over their participation status and data contributions using two main measures. First, we adopt an “opt-in” participation mechanism, where participants actively add collectors to their Crepe app to join data collections. We intend this to be combined with the explicit consents that researchers must obtain before each data collection study starts. Second, each participant can opt out of any collection, at any time, by deleting the collector from their Crepe app, which will immediately terminate the collection on their mobile devices.
For full transparency, the participant can in real-time check the examples of data being collected from their phone in the app’s collector details popup. Whenever a collector is modified by the researcher (e.g. extended or shortened collection period, additional or reduced types of data being collected), all participants will be informed of the changes through push notifications. We are also continuously rolling out new privacy-centric features, such as a data visualization dashboard for participants to see their contributed data volume and content, and the ability to request all their data be deleted when they withdraw from a study.
@@ 3.2.1. Building blocks of graph queries: screen entities and their relations
Figure 3. Details regarding the process of query generation and execution in Crepe. Section A (1) shows a UI Snapshot with UI elements and their relations, from which potential graph queries are generated (2) by combining screen entity relations to uniquely identify the target UI element on this UI Snapshot. Section B (3) depicts the execution of a chosen graph query on a different UI Snapshot, where the target element “duecelove” is successfully located, while the query fails to find a match on another screen.
@@
3.2.2. Automatic creation of graph queries
When a researcher specifies a piece of data to collect through the demonstration of tapping on the target data on screen (e.g. tapping on “apple” on the screenshot in Figure 3 section 1), we locate the associated UI element in the UI Snapshot, and generate a list of potential graph queries based on the target data’s relations with other screen entities in the UI Snapshot (2).

### v2

@@ Abstract.
Collecting mobile screen information datasets remains challenging for academic researchers. Commercial organizations often have exclusive access to mobile data, leading to a “data monopoly” that restricts academic research and user transparency. Existing open-source mobile data collection frameworks primarily focus on mobile sensing data rather than screen content. We present Crepe, a no-code Android app that enables researchers to collect information displayed on screen through simple demonstrations of target data. Crepe utilizes a novel Graph Query technique, which augments mobile UI structures to support flexible identification, location, and collection of specific data pieces. The tool emphasizes participants’ privacy and agency by providing full transparency over collected data and allowing easy opt-out. We designed and built Crepe for research purposes only and in scenarios where researchers obtain explicit consent from participants. Code for Crepe will be open-sourced to support future academic research data collection.
journalyear: 2024
@@ 1. Introduction
To conduct research on mobile information consumption and behavior, researchers often turn to screen data, the rich information displayed on users’ mobile interfaces, as a window into the information that algorithms show people and how people interact with technology in everyday lives. Researchers analyze mobile device screen info datasets to uncover insights in Human Computer Interaction (HCI), social sciences, and adjacent fields (Raento et al., 2009). Unlike traditional mobile sensing data that captures device states and sensor readings, screen information represents what users actually see and engage with, making it an invaluable resource for understanding the user side of interactions (Brown et al., 2013; Rogers and Marshall, 2017; Bentley et al., 2016).
However, collecting mobile screen data has remained frustratingly elusive for academic researchers. In practice, access to comprehensive mobile app usage data is often exclusively controlled by smartphone platforms like iOS and Android, or by individual app developers who guard their user data closely (Chandramouli et al., 2015; Generosi et al., 2020). Academic researchers find themselves in a challenging position: they must either rely on limited public developer APIs—which are subject to corporate policy changes and restrictions (Reddit, 2023; Barnes, 2023)—or negotiate complex collaborations with commercial organizations that may compromise research independence (Kramer et al., 2014). This “data monopoly” creates significant barriers to empirical mobile research (Jia-Jun Li et al., 2022).
@@
However, many researchers have limited resources and technical expertise to build their own data collectors. Meanwhile, the availability of developer APIs are subject to corporations’ discretionary policies (Reddit, 2023; Barnes, 2023). When researchers collaborate with commercial organizations to access mobile data (Kramer et al., 2014; Deka et al., 2017), they often need to align with the organizations’ business goals and public images, restricting the independence of academic research (Kramer et al., 2014). These phenomena are recognized as commercial organizations’ “data monopoly” in previous research (Jia-Jun Li et al., 2022) and pose challenges for empirical studies of mobile screen data.
@@
While numerous open-source mobile data collection frameworks exist to support academic research, most mature solutions focus on mobile sensing data rather than the actual screen information that users see and interact with (Ferreira et al., 2015; Aharony et al., 2011; Cardone et al., 2013). The technical challenge lies in reliably and automatically identifying which user interface screens contain target data, then locating and collecting that specific information. Current approaches often require users to actively input data or upload screenshots manually (Calacci and Pentland, 2022), or they continuously record user screens—an approach that is both inefficient and raises serious privacy concerns (Krieter, 2019).
In this work, we ask: how can we democratize mobile screen data collection for academic researchers, while respecting user privacy and maintaining data quality?
We created Crepe [^1] Crepe is an acronym for “Collector for Research Experiments of Participant Experiences”, a novel Android data collector app that embodies a programming-by-demonstration approach to mobile screen data collection. At its core, Crepe introduces Graph Query, a query language that augments mobile UI screen structures to support flexible identification, location, and collection of specific data pieces. With Crepe, researchers can define what data to collect through simple, no-code demonstrations, while participants maintain full transparency and control over their data through intuitive privacy controls and real-time collection feedback.
We designed Crepe with participant agency at its center. Users see exactly what data is being collected through semi-transparent overlays and dedicated data pages, and they can leave any study at any time by simply removing the collector from their app. This approach transforms mobile data collection from an opaque, researcher-controlled process into a transparent, participant-empowered collaboration.
To evaluate Crepe, we conducted a series of three user studies with both researchers and participants, uncovering performance characteristics, limitations, and future directions in both laboratory and real-world settings.
@@
    A novel query language, Graph Query, to reliably identify, locate, and collect target data on UI screens;
@@
    Crepe, the first Android data collector app to support flexible, customizable, low-code UI screen data collection using programming by demonstration (PBD) (Cypher and Halbert, 1993);
@@
    A series of three user studies demonstrating Crepe’s effectiveness in empowering both researchers and participants in mobile screen data collection.
@@ 2.1. Understanding Human Activities via Mobile Data
Mobile devices, especially smartphones, are now ubiquitous in everyday lives. Mobile applications harness various information about user social activities, which are embedded in users’ interactions, app usage patterns, and preferences. As a result, researchers in HCI, and social science have been leveraging rich data gained from mobile apps to better improve user experiences (Deka et al., 2017; Wang et al., 2023), uncover societal trends (Vieweg et al., 2010; Starbird et al., 2014; Starbird and Palen, 2010), and understanding user behaviors and attitudes (Fakhradyan, 2021; De et al., 2017; Balaban et al., 2022; Kramer et al., 2014).
@@
2.1.1. Analyzing Behavioral Data Streams
Researchers from diverse backgrounds, such as social science or business, mine mobile data for behavioral patterns and insights. Experience Sampling Method (ESM) is a research methodology to obtain in situ data for constructing an understanding of subjects’ daily behaviors, feelings, and thoughts of participants (van Berkel et al., 2017; Yue et al., 2014). In the survey conducted by Berkel et al., smartphone-based ESM benefits researchers in improving data quality through validation, collecting rich multimodal data for context reconstruction, and enabling real-time data collection and analysis (van Berkel et al., 2017). ESM can provide social scientists with combined human and sensor data in addition to background logging (Raento et al., 2009). Aside from sensor data, ESM questionnaires can also capture external aspects of experience, such as time, place, activities, and companion (Hektner et al., 2007). For instance, Yue et al. adopted ESM to study the function of photos. They found that photos can be used to trigger memory during follow-up interviews and as a beneficial component in data analysis (Yue et al., 2014). In this work, our goal is to create a low-code mobile UI screen data collection tool. Our app Crepe can facilitate academic researchers to instrument screen data collections without the need to extensively develop customized data collectors, or have public access to data APIs.
@@ 3. Proposed Method
Our goal for the Crepe data collector is three-fold. First, we seek to empower researchers lacking programming expertise to easily create and deploy data collections. Second, we prioritize the privacy and agency of individual data collection participants. Lastly, we also want to address the following observed technical challenges in reliable mobile screen data collection, which hinder even technical researchers:
@@
    The difficulty in automatically detecting that target data showed up on screen;
@@
    The lack of precision in locating target data and accessing its content;
@@
    The limited generalizability for dynamically changing data content.
To solve these challenges, previous collectors have asked participants to either directly input target data or upload screenshots or screen recordings containing target data (Calacci and Pentland, 2022; Krieter, 2019). However, these practices significantly increase participants’ efforts and reduce the reliability of the collected data’s quality.
In contrast, in Crepe, we used a novel Graph Query technique to address the above challenges. Graph Query has the following characteristics:
  1. (1)
    Triggering a data collection only when the target data shows up;
  2. (2)
    Locating and accessing only the target data on screen;
  3. (3)
    Generalizing easily to diverse data types, especially dynamically changing data.
@@ 3.1. The Crepe Data Collector User Experience
Figure 2. The workflow of Crepe for our main user groups: data collection researchers and participants. Researchers create a new data collector by demonstration A and share the collector ID with participants B. Participants add the collector to their own devices C, which runs in the background to collect the specified target data D. Note that in step D, Crepe uses Android Accessibility Service to access the view hierarchy of the current UI screen, then processes the view hierarchy, instead of directly working with screenshot images. The collected data is transmitted to a database E for the researcher to analyze and gain insights F. The colors in the background indicate each user group’s experience involved in the holistic Crepe pipeline.
@@
The figure illustrates the workflow of the Crepe system with 6 labeled steps: A) The researcher creates a new data collector by specifying the time period, app package, data field and description through a graphical interface on their device. B) The researcher shares the collector ID with participants. C) Participants add the collector to their own devices by entering the shared collector ID into the Instagram app running on their phone. The collector details like start/end time are shown. D) On the participant’s device, the Crepe app automatically identifies and locates the target UI data elements (like “walmart”, “deucelove”) using Graph Query, based on the UI screenshot. E) The target data elements collected by Crepe on participant devices get transmitted to a database. F) The researcher analyzes the collected data from the database to gain insights for their study. Arrows show the flow of information between the researcher device, participant devices, and the data collection database.
@@
Researchers have two main tasks when using Crepe (Figure 2): creating a collector A, and sharing collectors with study participants B. A researcher can use Crepe’s graphical user interfaces to define the start and end dates of a collector, as well as the target app to collect data from.
@@
The most important yet trickiest step in defining a data collector is specifying the target data on screen to collect. In Crepe, our design goal is to minimize the complexity of this specification process for researchers, by only asking them to tap on the screen data to collect in the target app, and describe their intention to collect these data (Figure 2 A). Crepe automatically translates these two pieces of information into a formal, executable Graph Query, which includes information about the target data’s parent app package, component type, and relations with other UI entities on screen. This Graph Query is the core of a data collector and will be used to reliably identify and precisely locate target data on the study participants’ devices. This specification process follows a programming-by-demonstration paradigm (Cypher and Halbert, 1993) and provides a simple, no-code experience for researchers to specify the target data. (The technical detail will be further explained in the following subsection on Graph Query.) After the target data are specified and descriptions are added, a collector is created. Afterward, the researcher can copy and share the unique ID of this collector with participants B. The collector will run on the participants’ device for the specified data collection periods. Afterwards, the researcher can receive the data collected in a self-defined database (E) and analyze it to uncover research insights (F). Our initial implementation of Crepe transmits collected data to a pre-configured Google Firebase Realtime Database that encrypts data both in transit and at rest [^3] https://firebase.google.com/support/privacy. We plan to support researchers’ customized configuration of their own database servers in the future.
@@
For study participants, they can easily add a collector to Crepe by entering the unique collector ID shared by the study’s researchers C, using Crepe’s graphical user interfaces. After granting Crepe with the necessary Android Accessibility Service permission, the added collector(s) will start automatically running in the background of the participants’ devices (D). Guided by the Graph Queries associated with the participants’ added collectors, Crepe will selectively seek the appearance of target data exclusively when the designated target apps are active. Once the target data is detected, Crepe will scrape and transmit the relevant information to the researchers (E). The participant will be able to monitor the active status of the collector(s) they have joined in Crepe. To drop out of a study, participants can simply delete a collector from their Crepe app, and the collector will no longer run in the background.
Figure 3. Two design features that enhance data collection transparency for system users (data contributors). When the system operates in the background, a yellow highlight appears over the collected content at each time of collection ( 1). Additionally, we designed and implemented a page that displays all collected data organized by collector, allowing users to review their contribution history and details ( 2).
[…]
This image displays two key transparency features of a data collection system interface: The left side shows multiple overlapping interface elements, including a map in the background, an Uber ride selection menu, a news article with a headline about cryptocurrency, and a Facebook promotional card. A yellow highlight surrounds the active content being collected, drawing attention to what data is being captured. This section is labeled “1 Visual highlight for collected data.” The right side shows a data dashboard interface titled ”My Data” that provides users with a detailed log of their collected data. The dashboard displays information for two data collectors: Instagram and Uber. For each collector, it shows when collection started (January 29, 2025 for Instagram, February 1, 2025 for Uber), the scheduled end date (February 8, 2025 for both), and the total number of data entries (46 for Instagram, 19 for Uber). The dashboard also displays individual data collection events with timestamps, event types, and content descriptions, including specific details about Uber rides such as fare information, pickup times, and driver proximity. This section is labeled “2 Data dashboard displaying all collected data.” Together, these features provide transparency by highlighting data as it’s collected and giving users access to a comprehensive record of their contributed data.
Data collection transparency
@@
Ensuring a transparent communication of data collection is critical in our design of Crepe. We ensure participants have the full control over their participation status and data contributions using two main measures. First, we adopt an “opt-in” participation mechanism, where participants actively add collectors to their Crepe app to join data collections. We intend this to be combined with the explicit consents that researchers must obtain before each data collection study starts. Second, each participant can opt out of any collection, at any time, by deleting the collector from their Crepe app, which will immediately terminate the collection on their mobile devices.
We also prioritized data collection transparency in Crepe design Figure 3. Every time data is collected from the user’s screen, a semi-transparent yellow highlight will appear over the collected data every time ( 1). We also designed a dedicated screen to show the users all of the data collected ( 2). We will also continuously release new privacy-centric features, such as daily or weekly push notifications summarizing data contributions, and the ability to request all their data be deleted when they withdraw from a study.
@@ 3.2.1. Building blocks of Graph Queries: screen entities and their relations
Figure 4. The detailed process of Graph Query generation and execution in Crepe. Section A shows how to generate a Graph Query: Crepe first uses Android Accessibility Service to access the view hierarchy of a UI screen 1 and augment it into a UI Snapshot graph, using UI elements’ characteristics and their relations. To identify the target UI element within the UI Snapshot graph, Crepe combines a set of UI element characteristics on this UI Snapshot to uniquely identify the target UI element on this UI Snapshot. We put these characteristics together to construct our Graph Query 2 and rank them based on their flexibility. In Section B, step 3 depicts the execution of a chosen Graph Query on two new screens, one containing the target data and one not. The set of unique characteristics of the target UI element ensures we only successfully collect the target data in “duecelove”, but not on the other screen.
@@
3.2.2. Automatic generation of Graph Queries
When a researcher demonstrate the target data to collect by tapping on the it on screen (e.g. tapping on “apple” on the screen in Figure 4 Step 1), we locate the associated UI element in the UI Snapshot, and utilize a unique set of its characteristics in the UI Snapshot to generate Graph Query candidates (Figure 4, Step 2).